# VENI, VINDy, VICI – a variational reduced-order modeling framework with uncertainty quantification

### v1

@@ 3.3.1 Results
where the cubic term, with $\gamma>0$, accounts for the hardening response and the absence of quadratic terms should be remarked. As the objective of VINDy is to identify a sparse reduced system, we wish to unveil the normal form underlying the observed high-dimensional system, directly from data. For the considered example, we can see from Fig. 11 that the terms indicated as relevant by our proposed method correspond to the ones present in the normal form (23) plus an additional bias term of small magnitude. Moreover the actual coefficients $\omega_{0}$ and $\xi$ could be retrieved from the spectral properties of the linearization of system (21). These coefficients represent the natural frequency and the damping of the device, respectively, and uniquely determine the linear terms in the reduced dynamics. In the considered problem, $\omega_{0}=0.5475\,\text{rad}/\mu\text{s}$ and $\xi=0.01\,{\mu\text{s}}^{-1}$. Our method successfully manages to accurately identify the linear terms $-\omega_{0}^{2}$ and $2\xi\omega_{0}$ (see Fig.11), from noisy data, without accessing or incorporating knowledge from the full order model.
Figure 11: Beam coefficient distribution. The terms that appear in the normal form of the observed high-dimensional system are framed in red. The cutoff value that determines which terms are kept is shown as dashed line. Moreover, the correct values that result from the natural frequency and the damping of the system are marked with a black triangle at the bottom of the according axis.
Figure 12: Physical reconstruction of latent mean trajectories and corresponding uncertainties for a node in the middle of the beam where the most dynamics occur for two example test simulations. For training only the first $T_{\text{train}}=1091\,\mu$s are used; the remaining time for which the model extrapolates has a red background.
Finally, we demonstrate the performance of the proposed method to generate the full-time evolution of the observed system for new parameter configurations following the VICI procedure. In Fig. 12, we illustrate that the reconstruction of the integrated latent dynamics accurately follows the reference. The trajectory resulting from the means of the coefficient distribution closely aligns with the actual course and the standard deviations resulting from the VICI approach cover the discrepancy throughout the trajectories.
To get a better visualization of the performance over the entire beam, we present in Fig. 13 a complete representation of the deformed beam at different points in time showing a contour plot of displacement errors. Clearly recognisable, the prediction follows the reference solution satisfactorily over the time period considered (including extrapolation in time).
@@
4 Discussion and concluding remarks
The paper introduces a novel framework – named VENI, VINDy, VICI – for creating interpretable, generative reduced-order models that can simulate high-dimensional, nonlinear dynamical systems with embedded uncertainty quantification. The proposed method efficiently provides accurate and reliable estimates by combining variational autoencoders with a newly introduced variational identification of nonlinear dynamics (VINDy) technique. VINDy identifies an interpretable, probabilistic model for the reduced dynamics by learning the distribution of coefficients of the linear combination of a parsimonious set of functions selected from a library of candidates.
As a result, in a purely data-driven and automated fashion, our overall method simultaneously learns the probability distribution of low-dimensional representation of the system states together with a probabilistic dynamical model, describing how the states evolve in time. Moreover, by adeptly accounting for uncertainty during both training and testing phases, the method operates effectively in high-noise and low-data regimes, and provides uncertainty-aware predictions for varying parameter configurations and initial conditions of the system.
The method’s effectiveness is showcased across various benchmarks, including the low-dimensional Rössler system to assess handling different sources and intensities of uncertainty, and high-dimensional PDE problems in fluid dynamics with a parametrised reaction-diffusion problem, and in structural mechanics with the hardening behavior of a micro electro-mechanical resonator.
A key feature of the method is the usage of variational inference to reformulate UQ as an optimization problem, solvable with efficient gradient-based techniques. This significantly reduces computational costs with respect to alternative strategies based on Bayesian or Monte Carlo methods. Moreover, it allows to seamlessly integrate UQ with the learning tasks into a single, training procedure.
As drawback, as with most deep learning frameworks, the proposed method can be sensitive to the weights initialization, choice of the priors and hyperparameters, such as, e.g., the weighting coefficients leveraging the contribution of the different loss terms in (17). In Sect. 2.4 we provided a general guidance on how to set these hyperparameters, but automatic, and more sophisticated hyperparameter-tuning techniques or constrained learning could complement the current framework.
A further improvement could be achieved by considering more advanced neural network models as building block of the variational autoencoders. For instance, graph neural networks [47] are more expressive in reconstructing spatial pattern so they can substitute the feed-forward neural networks. Indeed, while the method showcases accurate forecast of dynamics and extrapolation with respect to time thanks to the identified, explicit dynamical model, we do not expect as good generalization capabilities from feed-forward networks in predicting spatial patterns that are totally unseen over the training.
Finally we observe that, while UQ on the dynamical model could be inferred directly from the identified posterior distributions of the model coefficients, performing UQ over the predicted time trajectories, instead requires a sampling strategy, ı.e. VICI. Indeed, the reduced states are treated as independent and identically distributed (iid) variables, making them completely agnostic of time. So UQ over trajectories is computed by VICI from the statistical moments of different integrated trajectories from different realizations of initial conditions and coefficients of the dynamical model. Further developments could model latent states as random processes, thus accounting for time dependence and providing more rigorous UQ for the predicted trajectories.
The non-intrusive, data-driven nature of the framework requires minimal a priori knowledge of the observed system and could handle model and measurement noise. This paves the way for applications to real-life datasets, potentially considering various data types, such as those coming from sensors or videos as in [27, 7]. As further direction of future work, the flexibility and generality of the framework offers promising prospects for its extension to online learning within the context of digital twins, enabling dynamic model updates as new data emerge. This could be done by adopting previous posterior estimates as priors and fine-tuning the model to refine posterior distributions.
@@ Acknowledgment
The source code and data for the VENI, VINDy, VICI Framework, examples, and simulation setups are available at https://github.com/jkneifl/VENI-VINDy-VICI. All code is implemented in Python using TensorFlow.
Appendix A Appendix
@@ A.1 VINDy Loss
In this section we derive equation (13) in Section 2.2. We assume that $\bm{\Xi}$ and $\bm{z}$ are independent. This reflects the fact that $\bm{\Xi}$ are constant coefficients determining the contribution of each candidate feature in the dynamics, thus they are related to the underlying physical model we aim to discover and not to the specific (latent) observed data.
In VINDy, we approximate the distribution of $p(\bm{\Xi}|\dot{\bm{z}})$ of $\bm{\Xi}|\dot{\bm{z}}$ with a posterior distribution $q(\bm{\Xi})$ parameterized by trainable weights. We aim to minimize the KL divergence between the approximated posterior and the true one:
$\begin{split}\text{KL}(q(\bm{\Xi})\parallel p(\bm{\Xi}\mid\dot{\bm{z}}))&=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log q(\bm{\Xi})]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log p(\bm{\Xi}\mid\dot{\bm{z}})]\\ &=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log q(\bm{\Xi})]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log p(\bm{\Xi}\mid\dot{\bm{z}},\,\bm{z})]\\ &=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log q(\bm{\Xi})]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log\frac{{p(\bm{\Xi},\,\bm{z})}p(\dot{\bm{z}}\mid\bm{\Xi},\,\bm{z})}{p(\dot{\bm{z}}\mid{\bm{z}})p({\bm{z}})}\right]\\ &=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log q(\bm{\Xi})]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log\frac{p(\bm{\Xi})p(\bm{z})p(\dot{\bm{z}}\mid\bm{\Xi},\,\bm{z})}{p(\dot{\bm{z}}\mid{\bm{z}})p({\bm{z}})}\right]\\ &=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log q(\bm{\Xi})]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log p(\bm{\Xi})\right]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log(p(\dot{\bm{z}}\mid\bm{\Xi},\,\bm{z}))\right]+\log p(\dot{\bm{z}}\mid{\bm{z}})\\ &=\text{KL}(q(\bm{\Xi})\parallel p(\bm{\Xi}))-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log(p(\dot{\bm{z}}\mid\bm{\Xi},\,\bm{z}))\right]+\log p(\dot{\bm{z}}\mid{\bm{z}}).\end{split}$ (24)
@@
Analogously, the KL divergence between $p(\bm{z}|\dot{\bm{z}})$ and its approximation by means of the posterior $q(\bm{z})$ can be written as
@@
$\begin{split}\log p(\dot{\bm{z}})&-\text{KL}(q(\bm{\Xi})\parallel p(\bm{\Xi}|\dot{\bm{z}}))-\text{KL}(q(\bm{z})\parallel p(\bm{z}|\dot{\bm{z}}))=\\ &=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log(p(\dot{\bm{z}}\mid\bm{\Xi},\bm{z}))\right]-\text{KL}(q(\bm{\Xi})\parallel p(\bm{\Xi}))-\text{KL}(q(\bm{z})\parallel p(\bm{z})),\end{split}$
@@ A.2 Prior Distributions
The priors $p(\bm{\Xi})$ and $p(\bm{z})$ in the context of VENI and VINDy can be used to infuse knowledge about the distributions of the coefficients of the dynamics and latent states. If no prior knowledge about the distribution’s shape and its parameters is present, Gaussian priors are a common choice in variational inference, especially in VAEs. Moreover, using latent Gaussian distributions for the latent variables allows modeling arbitrarily complex output distributions if the decoder is sufficiently complex [48]. A valid alternative, especially when model the dynamics’ coefficients is the use of Laplacian priors. In contrast to Gaussian priors, Laplacian priors can act as sparsity-promoting regularization terms within a regression task [26]. Consequently, they are more prone to provide a parsimonious dynamical model, which is in general highly desirable [49]. We furthermore want highlight that adjusting the priors are an elegant way of incorporating any existing pre-knowledge about occurring function terms. Other possible choices for the priors may include spike and slab, regularized horseshoe, etc. In this work, we focus on Gaussian and Laplacian priors, since they have the additional advantage that the KL divergence between two distribution of the same family can be expressed in a closed-form, as presented in the following.
@@
is parameterized by its mean ${\mu}^{\mathcal{N}}$ and variance ${{\delta}^{\mathcal{N}}}$. For two univariate Gaussian distributions the KL divergence can be calculated using the closed form
$\text{KL}(\mathcal{N}({\mu}^{\mathcal{N}}_{1},{\delta}^{\mathcal{N}}_{1})\parallel\mathcal{N}({\mu}^{\mathcal{N}}_{2},{\delta}^{\mathcal{N}}_{2}))=\log(\frac{{\delta}^{\mathcal{N}}_{2}}{{\delta}^{\mathcal{N}}_{1}})+\frac{{{\delta}^{\mathcal{N}}_{1}}^{2}+({\mu}^{\mathcal{N}}_{2}-{\mu}^{\mathcal{N}}_{1})^{2}}{2{{\delta}^{\mathcal{N}}_{2}}^{2}}-\frac{1}{2}.$ (28)
@@ A Laplace distribution
is parameterized by a location ${\mu}^{\mathcal{L}}$ and a scale parameter ${\delta}^{\mathcal{L}}$. For gradient-based optimization the reparameterization trick for a Laplace distribution looks similar as the one for Gaussian ones
@@
$\text{KL}(\mathcal{L}({\mu}^{\mathcal{L}}_{1},{\delta}^{\mathcal{L}}_{1})\parallel\mathcal{L}({\mu}^{\mathcal{L}}_{2},{\delta}^{\mathcal{L}}_{2}))=\log(\frac{{\delta}^{\mathcal{L}}_{2}}{{\delta}^{\mathcal{L}}_{1}})+\frac{{\delta}^{\mathcal{L}}_{1}\exp{-\frac{|{\mu}^{\mathcal{L}}_{1}-{\mu}^{\mathcal{L}}_{2}|}{{\delta}^{\mathcal{L}}_{1}}}+|{\mu}^{\mathcal{L}}_{1}-{\mu}^{\mathcal{L}}_{2}|}{{\delta}^{\mathcal{L}}_{2}}-1,$ (31)

### v2

@@ Results
where $\omega_{0}^{2}$ represents the natural frequency and $2\xi\omega_{0}$ the damping of the device. Moreover, the cubic term, with $\gamma>0$, accounts for the hardening response and the absence of quadratic terms should be remarked.
@@
As the objective of VINDy is to identify a sparse reduced system, we wish to unveil the underlying normal form (22) of the observed high-dimensional system, directly from data. For the considered example, we can see from Fig. 5 that the terms indicated as relevant by our method correspond to those present in the normal form plus an additional bias term of small magnitude. Our method effectively and accurately identifies the actual coefficients of the linear terms associated with the natural frequency and the damping, see Fig.5. This is achieved without requiring access to or integration of knowledge from the FOM, relying solely on noisy data. In addition, the full dynamics term in the loss function (16) promotes consistency between the latent dynamics and the decoded trajectories, allowing key nonlinear features of the MEMS system – such as the hardening response, loss of uniqueness, and jump phenomena – to be preserved after decoding [13]. Fig. 5 also shows that the mean full-time evolutions produced by the VICI procedure for new parameter configurations follow the reference, while the standard deviations capture the discrepancies.
To assess the performance of the proposed method, we provide in B a benchmark comparison against state-of-the-art approaches, including Fourier neural operators [45] and Gaussian Process-based ROMs, highlighting the unique features of our framework and main advantages in terms of accuracy, computational efficiency, scalability, uncertainty quantification, extrapolation capabilities and physical interpretability.
4 Discussion and final remarks
The results demonstrate the effectiveness of the proposed VENI, VINDy, VICI framework in constructing accurate, interpretable, and uncertainty-aware ROMs directly from time-series data. This is achieved by integrating probabilistic modeling with data-driven discovery in a unified physical generative framework. The identified ROMs are consistent with the underlying physics, preserving computational efficiency while providing robust UQ, even in situations with significant noise or limited data. VINDy’s capability to select essential terms in the dynamics allows for the discovery of governing equations that offer insights into the fundamental mechanisms of complex systems. In B, we present a benchmark comparison that highlights the unique characteristics of the proposed framework and its advantages relative to alternative state-of-the-art modeling approaches, including Fourier neural operators [45] and Gaussian process-based ROMs.
By modeling both state variables and model coefficients as probability distributions, our framework accommodates measurement noise and model uncertainty, enhancing the robustness and reliability of the ROM’s predictions. Several advantages are noted and exploited. First, the estimated distributions of model coefficients provide valuable insights into the uncertainties in function terms, guiding model selection and refinement. Second, sparsity-promoting priors, such as Laplacian priors, in combination with PDF thresholding provide parsimonious models [26]. Sparsity is fundamental for interpretability and generalization, as it enables the identification of essential terms in the dynamics, thereby revealing the governing equations underlying the observed system among an infinite range of possible models. For example, in the Rössler system, our method accurately identifies the governing equations by estimating the essential terms and their coefficients. In the reaction-diffusion system, the primary dynamics is effectively modeled as a linear oscillator, with the phase space in the latent coordinates accurately approximating the attracting limit cycle. In the MEMS resonator application, the framework autonomously discovers the governing dynamics, unveiling the normal form of the system. This representation not only captures the dynamics in the sparsest form but also provides direct insights into the linear coefficients (e.g., natural frequency and damping) and nonlinear terms determining the system’s dynamics.
Unlike traditional methods that enforce sparsity based solely on coefficient magnitude, the proposed framework leverages probability density functions of the coefficients. This allows coefficients with small magnitudes but substantial impacts on the system’s dynamics to be retained. For example, in the beam case, the forcing term – despite its small amplitude – is essential for accurately capturing the system’s non-autonomous behavior. While magnitude-based approaches might discard this term, our method retains it, highlighting its importance.
In addition to the UQ directly provided by the approximated posterior on latent states and dynamical models, the framework enables UQ on physical solution trajectories using the ensemble strategy of VINDy. The generative nature of the method allows the creation of an ensemble of dynamical models through sampling, at virtually no additional cost, unlike standard ensemble-based UQ methods that require training of multiple instances of a model. By integrating these models over time and mapping the resulting ensemble of latent trajectories to physical space, we gain direct insights into how uncertainty in the spatial solutions evolves over time.
While uncertainties are currently modeled globally at the system level – reflecting the assumption of a single set of time-invariant governing equations with parametric dependence – the framework admits a natural extension to component-wise structure by assigning dedicated latent variables and corresponding governing equations to distinct subsystems, thereby localizing uncertainty across system partitions when warranted by the application.
As with most deep learning approaches, the performance of our method can be sensitive to hyperparameters, including the weighting of loss terms, the choice of priors, and the VINDy library of candidate functions. Another limitation is that, while the approach effectively forecasts solution trajectories via the identified latent dynamical model, the decoder does not guarantee strong generalization to spatial patterns entirely outside the training data, which may result in unphysical reconstructions in extrapolation scenarios. This limitation could be alleviated by employing more advanced architectures as building blocks for the VAE; for instance, graph neural networks [46] may improve spatial pattern reconstruction and enhance generalization to unseen states. Moreover, a systematic analysis of reconstruction errors and potential physical violations under varying decoder performance represents an important direction for future work.
While the framework demonstrates broad applicability, its performance depends on several modeling assumptions. In particular, an adequate low-dimensional latent representation is crucial. If the system requires higher-dimensional embeddings, the increasing number of interaction terms may affect robustness and interpretability. Similarly, very high measurement noise can challenge inference and dynamics reconstruction, although the correct model terms may still be identified. The approach also builds on the assumption that latent dynamics can be captured in a sparse form; when this is not the case, solutions may be less interpretable or less efficient. Finally, a misspecified prior can bias the inference process and yield inaccurate uncertainty estimates.
Although the framework demonstrated robustness to noise, simulation-generated datasets cannot fully reflect the uncertainties, sensor imperfections, and modeling discrepancies inherent in real-world experiments. Consequently, dealing with more realistic measurement artifacts (e.g., impulsive, drift, or correlated noise) and applying the method to real-life datasets, including sensor and video data [47, 11], represent exciting future directions to validate its practical robustness – particularly given its non-intrusive nature. As a further direction of development, the framework’s flexibility offers strong potential for online learning, allowing dynamic model updates as new data emerge. This could be achieved by using previous posterior estimates as priors and fine-tuning the model to refine posterior distributions, potentially in combination with conformal prediction techniques [48, 49] to provide theoretically grounded uncertainty intervals under distribution shift and non-stationarity.
@@ Acknowledgment
The source code and data for the VENI, VINDy, VICI Framework, examples, and simulation setups are available at https://github.com/jkneifl/VENI-VINDy-VICI with a persistent release in [50]. All code is implemented in Python using TensorFlow.
Appendix A VENI–VINDy–VICI: methodological and implementation details
@@ A.1 VINDy Loss
In this section we derive VINDy objective function (12) in Sect. 2.3. We assume that $\bm{\Xi}$ and $\bm{z}$ are independent random variables. This reflects the fact that $\bm{\Xi}$ are constant coefficients determining the contribution of each candidate feature in the dynamics, thus they are related to the underlying physical model we aim to discover and not to the specific (latent) observed data.
In VINDy, the posterior distribution $p(\bm{\Xi}|\dot{\bm{z}})$ is approximated with $q(\bm{\Xi})$ parameterized by trainable weights. For the aforementioned reason, the approximated posterior for $\bm{\Xi}$ does not feature an explicit dependency on the conditioning variable $\dot{\bm{z}}$. The weights parameterizing $q$ are optimized by minimizing the KL divergence between the approximated posterior and the true one:
$\begin{split}\text{KL}(q(\bm{\Xi})&\parallel p(\bm{\Xi}\mid\dot{\bm{z}}))=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log q(\bm{\Xi})]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log p(\bm{\Xi}\mid\dot{\bm{z}})]\\ &=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log q(\bm{\Xi})]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log p(\bm{\Xi}\mid\dot{\bm{z}},\,\bm{z})]\\ &=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log q(\bm{\Xi})]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log\frac{{p(\bm{\Xi},\,\bm{z})}p(\dot{\bm{z}}\mid\bm{\Xi},\,\bm{z})}{p(\dot{\bm{z}}\mid{\bm{z}})p({\bm{z}})}\right]\\ &=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log q(\bm{\Xi})]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log\frac{p(\bm{\Xi})p(\bm{z})p(\dot{\bm{z}}\mid\bm{\Xi},\,\bm{z})}{p(\dot{\bm{z}}\mid{\bm{z}})p({\bm{z}})}\right]\\ &=\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}[\log q(\bm{\Xi})]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log p(\bm{\Xi})\right]-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log(p(\dot{\bm{z}}\mid\bm{\Xi},\,\bm{z}))\right]+\log p(\dot{\bm{z}}\mid{\bm{z}})\\ &=\text{KL}(q(\bm{\Xi})\parallel p(\bm{\Xi}))-\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log(p(\dot{\bm{z}}\mid\bm{\Xi},\,\bm{z}))\right]+\log p(\dot{\bm{z}}\mid{\bm{z}}).\end{split}$ (23)
@@
Analogously, the KL divergence between $p(\bm{z}|\dot{\bm{z}})$ and the approximated posterior $q(\bm{z})$ can be written as
@@
$\begin{split}\log p(\dot{\bm{z}})&-\text{KL}(q(\bm{\Xi})\parallel p(\bm{\Xi}|\dot{\bm{z}}))-\text{KL}(q(\bm{z})\parallel p(\bm{z}|\dot{\bm{z}})){\approx}\\ &{\approx}\mathbb{E}_{\bm{\Xi}\sim q(\bm{\Xi})}\left[\log(p(\dot{\bm{z}}\mid\bm{\Xi},\bm{z}))\right]-\text{KL}(q(\bm{\Xi})\parallel p(\bm{\Xi}))-\text{KL}(q(\bm{z})\parallel p(\bm{z})),\end{split}$
@@ A.2 Prior Distributions
The priors $p(\bm{\Xi})$ and $p(\bm{z})$ in the context of VENI and VINDy can be used to infuse knowledge about the distributions of the coefficients of the dynamics and latent states. If no prior knowledge about the distribution’s shape and its parameters is present, Gaussian priors are a common choice in variational inference, especially in VAEs. Moreover, using latent Gaussian distributions for the latent variables allows modeling arbitrarily complex output distributions if the decoder is sufficiently complex [51]. A valid alternative, especially when modeling the dynamics’ coefficients is the use of Laplacian priors. In contrast to Gaussian priors, Laplacian priors can act as sparsity-promoting regularization terms within a regression task [26]. Consequently, they are more prone to provide a parsimonious dynamical model, which is in general highly desirable [17]. We furthermore want to highlight that adjusting the priors are an elegant way of incorporating any existing pre-knowledge about occurring function terms. Other possible choices for the priors may include spike and slab, regularized horseshoe, etc. In this work, we focus on Gaussian and Laplacian priors, since they have the additional advantage that the KL divergence between two distribution of the same family can be expressed in closed form, as presented in the following.
@@
the Gaussian distribution with mean ${\mu}^{\mathcal{N}}$ and variance ${\delta}^{\mathcal{N}}$. For two univariate Gaussian distributions, the KL divergence can be calculated in closed form as
$\text{KL}(\mathcal{N}({\mu}^{\mathcal{N}}_{1},{\delta}^{\mathcal{N}}_{1})\parallel\mathcal{N}({\mu}^{\mathcal{N}}_{2},{\delta}^{\mathcal{N}}_{2}))$ $=\log\left(\frac{{\delta}^{\mathcal{N}}_{2}}{{\delta}^{\mathcal{N}}_{1}}\right)+\frac{{{\delta}^{\mathcal{N}}_{1}}^{2}+({\mu}^{\mathcal{N}}_{2}-{\mu}^{\mathcal{N}}_{1})^{2}}{2{{\delta}^{\mathcal{N}}_{2}}^{2}}-\tfrac{1}{2}.$ (27)
@@
the Laplace distribution with location parameter $\mu$$^{\mathcal{L}}$ and scale parameter $\delta$$^{\mathcal{L}}$. For gradient-based optimization the reparameterization trick for a Laplace distribution looks similar as the one for Gaussian distributions.
@@
[…]